OpenStudy (anonymous):

Suppose that the attraction between the moon and the earth were due to Coulomb forces rather than gravitational force. What would be the magnitude of the charge required if equal but opposite charges resided on both earth and moon? Mass of earth = 5.98×1024 kg; mass of moon = 7.35×1022 kg; earth-moon distance = 3.84×108 m.

OpenStudy (anonymous):

Since the distance is the same, we just need that \[GMm = kq^2 \rightarrow q = \sqrt{\frac{GMm}{k}} \]

OpenStudy (anonymous):

thank you!